OpenStudy (anonymous):

Solve the system of equations using the elimination method. 3x-4y=0 2x-3y=-1

OpenStudy (anonymous):

how do i solve this?

OpenStudy (anonymous):

multiply the first equation by 2 and the 2nd equation by -3 and add them together

OpenStudy (anonymous):

2(3x-4y=0) => 6x-8y=0 -3(2x-3y=-1)=>-6x+9y=3 adding them together give you y=3

OpenStudy (anonymous):

6x-8y=0-eqn1 6x-9y=-3-eqn2 eqn1-eqn2=>y=3 6x-24=0 6x=24 x=4

OpenStudy (anonymous):

okay. got iy.

OpenStudy (campbell_st):

Label the equations 3x - 4y = 0 (1) and 2x - 3y = -1 (2) the 2 x (1) - 3 x (2) this will eliminate x as shown below 6x - 8y = 0 -6x+ 9y = -3 y = -3 hence y = -3 and sub into either equation to find x, x = 4

OpenStudy (campbell_st):

oops y = 3

OpenStudy (anonymous):

okay. thanks you guys! i really appreciate it! :)